The length of the rectangle exceeds its breadth by 3 cm. If the length and breadth are each increased by 2 cm, then the area of new rectangle will be 70 sq. cm more than that of the given rectangle. Find the length and breadth of the given rectangle.
Answer by ikleyn(52780)

.
One equation is 

L - W = 3     (1)        ("The length of the rectangle exceeds its breadth by 3 cm.")


The second equation is 

(L+2)*(W+2) - LW = 70,   ("If the length and breadth are each increased by 2 cm, then the area of new rectangle . . . ")  

or, after simplifying

L + W = 33     (2) 


So, you have the system of two equations in two unknowns (1),(2). By solving it,  you get

L = 18 cm,  W = 15 cm.


Check.  (18+2)*(15+2) - 18*15 = 70 sq cm.   ! Correct !

Solved.

------------------
To see other similar solved problems, see the lessons
    - Problems on the area and the dimensions of a rectangle surrounded by a strip
    - Problems on a circular pool and a walkway around it
in this site.